OpenStudy (anonymous):

solve log11 (y+8) + log11 4 = log11 60

OpenStudy (anonymous):

@mathmale

OpenStudy (whpalmer4):

\[\log_{11} (y+8) + \log_{11} (4) = \log_{11} (60)\] First thing to remember is that \[\log_a b + \log_a c = \log_a (b*c)\]

OpenStudy (whpalmer4):

If you apply that, what does the left hand side become?

OpenStudy (anonymous):

a is 11 b is y + 8 and c is 4 correct ?

OpenStudy (whpalmer4):

sure, if you want. The general notion is that the sum of the logs of two quantities is the log of the product of the two quantities.

OpenStudy (anonymous):

\[\log_{11} (y+8)(4)\]

OpenStudy (whpalmer4):

To be clear, \[\log_{11} (4(y+8))\] Now the right hand side is unchanged, so we have \[\log_{11} (4(y+8)) = \log_{11} (60)\]What does that imply about the relationship between \(4(y+8)\) and \(60\)?

OpenStudy (anonymous):

im not sure?

OpenStudy (whpalmer4):

Doesn't it mean they have to be equal? Doesn't \[\log x = \log y\] mean that \(x = y\)?

OpenStudy (anonymous):

oh. yea

OpenStudy (anonymous):

so now we find x and y?

OpenStudy (whpalmer4):

\[4(y+8) = 60\]

OpenStudy (anonymous):

solve for y

OpenStudy (whpalmer4):

yes, please.

OpenStudy (anonymous):

ok x = 8

OpenStudy (whpalmer4):

try again please :-)

OpenStudy (anonymous):

6 ***

OpenStudy (whpalmer4):

we want to find \(y\) such that \(4(y+8) = 60\) No other letters apply in this problem, the use of \(x\) was just as a placeholder in the example...it could have as easily been \[\log c = \log d\]

OpenStudy (anonymous):

ok so does x = 6?

OpenStudy (whpalmer4):

there is no \(x\) in this problem!

OpenStudy (anonymous):

so what is y ?

OpenStudy (whpalmer4):

and whatever you think your next answer is, try it in \(4(y+8) = 60\) before you post it.

OpenStudy (whpalmer4):

Solve the equation \[4(y+8) = 60\]for \(y\). What do you get?

OpenStudy (anonymous):

y = 7

OpenStudy (whpalmer4):

Yes. That is the answer.

OpenStudy (anonymous):

thanks!!

OpenStudy (whpalmer4):

Any questions about how we got there?

OpenStudy (anonymous):

just a little confused about how we set that equation up

OpenStudy (whpalmer4):

Which part, converting the addition of logs into the log of a product, or solving from there on?

OpenStudy (anonymous):

converting the addition

OpenStudy (anonymous):

It is 7

OpenStudy (anonymous):

4(y + 8) = 60 4y + 32 = 60 4y = 28 y = 7

OpenStudy (anonymous):

so, log11 (y + 8) + log11 4 = log11 60 it can rewriten be log11 (y + 8)4 = log11 60 now, cancel out the log 11 to both sides, giving us (y+8)4 = 60 solve for y so it will equal 7

OpenStudy (whpalmer4):

Well, let's back up a bit. Do you understand that a logarithm is just an exponent?

OpenStudy (anonymous):

i think thats correct

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Okay, so let's make it easy for calculating and use 2 as the base of our logarithms for this example. Let's say we have \[\log_2 4 + \log_2 8\]The logarithm base 2 of 4 is the number that we raise the base to get 4. 2^2 = 4. The logarithm base 2 of 8 is the number that we raise the base to get 8. 2^3 = 8. Okay so far?

OpenStudy (anonymous):

yes..

OpenStudy (whpalmer4):

Okay, let's add those two together: \(2+3 = 5\) Now, say we have a number \(x\) whose logarithm base 2 happens to be \(5\). Can you find the value of \(x\)? \[\log_2 x = 5\]\[x=\]

OpenStudy (whpalmer4):

Remember, if \[\log_2 x = 5\]that implies that \[2^5 = x\]

OpenStudy (anonymous):

yes i understand that

OpenStudy (whpalmer4):

Okay, so \(x =\)

OpenStudy (anonymous):

32

OpenStudy (whpalmer4):

Right. What does \(4*8=\)

OpenStudy (anonymous):

32

OpenStudy (whpalmer4):

Right. So \[\log_2 4 + \log_2 8 = 2+ 3 = 5\]and\[\log_2 32 = \log_2 (4*8) = \log_2 4 + \log_2 8 = 5\]This is no accident!

OpenStudy (whpalmer4):

\[b^m*b^n = b^{m+n}\] That's what we are doing when we add logarithms of a common base.

OpenStudy (anonymous):

ahh

OpenStudy (whpalmer4):

In our example, \[2^m*2^n = 2^2*2^3 = 2^{2+3} = 2^5\]

OpenStudy (whpalmer4):

Hopefully, it is also clear that subtracting logarithms is like dividing.

OpenStudy (anonymous):

i see now :D thanks so much!

OpenStudy (whpalmer4):

This is how old-school slide rules work — they have scales with logarithmic graduations, and the mechanics of the slide rule are just adding and subtracting lengths.

OpenStudy (whpalmer4):

In the old days, people used to publish books of tables of logarithms to a handful of decimal places. If you wanted to compute \(3.972^{4.026}\), you would look up the logarithm of \(3.972 \approx 0.599009\), multiply it by the exponent to get \(2.41161\), and then look that up in the table that goes from logarithms back to numbers to get \(257.995\).

OpenStudy (anonymous):

i see

OpenStudy (whpalmer4):

This also leads to a bad math joke: Noah's boat finally comes to rest as the flood waters recede, and he lowers the gangway and send the animals out calling to them, "Go forth and multiply". Most of the animals leave, but two snakes are left behind. Noah looks at them, and commands "Go forth and multiply!" The snakes look at him but do not move. He tries again, "Go forth and multiply!" The snakes do not move. Noah gets angry and in his most commanding voice shouts, "Go forth and multiply!" The snakes look up at him and say, "We can't, we're adders". Noah thinks for a while, then grabs his saw and hammer and runs off into the forest, where he cuts down a tree. He saws and hammers and builds a small table. He carefully picks up the snakes and puts them on the table. "Go forth and multiply!" he commands. The snakes look at each other, and then at Noah. "We can't, we're adders". "Yes", Noah replies, "but, even adders can multiply on a log table".

OpenStudy (anonymous):

LOL i don't get it but i will eventually

OpenStudy (whpalmer4):

Well, the key is that if you have log tables, you can multiply by adding. Add the logs of the terms in the product, take the anti-log to find the answer.